# SensorChat: Answering Qualitative and Quantitative Questions during Long-Term Multimodal Sensor Interactions

### v1

@@ Abstract.
Natural language interaction with sensing systems is crucial for enabling all users to comprehend sensor data and its impact on their everyday lives. However, existing systems, which typically operate in a Question Answering (QA) manner, are significantly limited in terms of the duration and complexity of sensor data they can handle.
In this work, we introduce SensorChat, the first end-to-end QA system designed for long-term sensor monitoring with multimodal and high-dimensional data including time series. SensorChat effectively answers both qualitative (requiring high-level reasoning) and quantitative (requiring accurate responses derived from sensor data) questions in real-world scenarios. To achieve this, SensorChat uses an innovative three-stage pipeline that includes question decomposition, sensor data query, and answer assembly. The first and third stages leverage Large Language Models (LLMs) for intuitive human interactions and to guide the sensor data query process. Unlike existing multimodal LLMs, SensorChat incorporates an explicit query stage to precisely extract factual information from long-duration sensor data. We implement SensorChat and demonstrate its capability for real-time interactions on a cloud server while also being able to run entirely on edge platforms after quantization. Comprehensive QA evaluations show that SensorChat achieves up to 26% higher answer accuracy than state-of-the-art systems on quantitative questions. Additionally, a user study with eight volunteers highlights SensorChat’s effectiveness in handling qualitative and open-ended questions.
@@ 1. Introduction
In recent years, the number of Internet of Things (IoT) devices in daily life has grown rapidly, including smartphones, smart home accessories, and smart city infrastructures (Choi et al., 2021; Khang et al., 2023). These devices generate a massive amount of multimodal sensor data, with total production expected to reach 80 ZB by 2025 (International Data Corporation (2021), IDC). Traditional research has focused on developing classification models for tasks such as human activity recognition (Ouyang et al., 2022, 2023; Xu et al., 2023b, a; Cao et al., 2024; Weng et al., 2024). In contrast, latest studies have recognized the importance of natural language interactions, especially for making sensor data more accessible and useful to elderly people or patients with Alzheimer’s (Xing et al., 2021; Nie et al., 2022b; Yang et al., 2024b; Arakawa et al., 2024; Ji et al., 2024; Chen et al., 2024). These systems typically operate in a Question Answering (QA) framework: users ask questions, the sensor system processes both the query and the underlying sensor data, and finally delivers an adequate answer. The questions can be qualitative, requiring high-level reasoning about a user’s lifestyle or well-being (e.g., “How is my work-life balance?”), or quantitative, demanding quantitatively precise answers based on sensor data (e.g., “How long did I walk?”). QA offers a natural and flexible way for users to interact with and gain insights from sensors, going beyond simple classification models.
(a) Example QA in Sensor2Text (Chen et al., 2024).
(b) Example QA in DrHouse (Yang et al., 2024b).
(c) Example quantitative (left) and qualitative (right) QA in SensorChat (our system).
Figure 1. A visual comparison of QA in state-of-the-art sensor systems and SensorChat. SensorChat excels in processing long-duration sensor data and generating quantitatively accurate answers and insightful qualitative analyses of well-being.
However, existing QA systems for sensors face significant limitations, particularly with short time windows or low-dimensional sensors. Most current systems can only process data over short periods, such as one minute in DeepSQA (Xing et al., 2021) or 16 seconds in Sensor2Text (Chen et al., 2024), as illustrated in Fig. 1(a). While useful for both qualitative and quantitative activity analysis, such short time frames offer limited value for gaining deeper insights into users’ fitness and overall well-being. Other QA systems focus on longer-term clinical diagnosis or health-related queries, but leverage low-dimensional physiological data like average heart rate or daily step counts (Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024b), as shown in Fig. 1(b). While effective for certain use cases, they cannot handle high-dimensional time series data, such as IMU signals, and fail to provide accurate answers about fine-grained activities. In practice, reasoning based on fine-grained physical activities, such as daily time spent exercising versus sitting, is critical for assessing a user’s health status (of Health et al., 2021). A significant gap remains in developing a QA system capable of answering both qualitative and quantitative questions using (1) long-duration and (2) high-dimensional multimodal sensor data.
@@
In this paper, we introduce SensorChat, the first end-to-end system designed for answering both qualitative and quantitative queries in long-term sensor monitoring, using multimodal and high-dimensional time series sensors. We consider the commonly used mobile devices for sensor data sources, i.e., smartphones and smartwatches. In contrast to prior QA systems (Xing et al., 2021; Nie et al., 2022b; Chen et al., 2024; Arakawa et al., 2024; Yang et al., 2024b), SensorChat is capable of processing long-duration sensor data spanning weeks or even months. For example, Fig. 1(c) shows a user’s activities over a three-day period, with the x-axis representing wall-clock time from 00:00 AM to 11:59 PM, and the y-axis denoting separate days. Different activity labels are depicted by color-coded bars. When asked a quantitative query like, ”What day did I spend the longest eating last week?” SensorChat provides a precise answer based on a detailed search of the sensor database, such as, ”You spent the most time eating on Saturday.” Furthermore, leveraging its detailed analysis of daily routines, SensorChat can also properly address open-ended qualitative queries related to overall health and lifestyle, such as, ”How would you rate my lifestyle?” To the best of the authors’ knowledge, SensorChat is the first QA system capable of processing such long-duration sensor data to provide both precise activity-based responses and insightful evaluations of lifestyle trends.
To achieve the goal, SensorChat introduces a novel three-stage design including question decomposition, sensor data query and answer assembly. The question decomposition and answer assembly stages utilize the power of Large Language Models (LLMs) to correctly understand arbitrary questions and provide accurate answers. SensorChat further enhances the LLMs in these stages via in-context learning and finetuning techniques to adapt LLMs for various QA scenarios. The intermediate sensor data query stage is pivotal in encoding raw time series data from multimodal sensors into a pre-stored embedding database. During offline pretraining, SensorChat uses a novel contrastive sensor-text pretraining loss to align sensor and text encoders. In the online stage, SensorChat accurately and exhaustively retrieves all relevant sensor embeddings by performing a similarity search against text embeddings generated from question decomposition. These queried embeddings are then processed by a set of carefully designed summarization functions to generate the sensor context in text. Finally, in the answer assembly stage, the sensor context is integrated with the original question using another LLM, to generate the ultimate answer. The explicit query mechanism and the fusion of textual sensor information with user questions distinguish SensorChat from emerging multimodal LLMs, which often rely on the ”black-box” integration of natural language and sensory data (Moon et al., 2023b; Zhang et al., 2023; Liu et al., 2024a; Han et al., 2024). As a result, SensorChat excels in answering long-term qualitative and quantitative questions accurately, where state-of-the-art QA systems fall short (Xing et al., 2021; Nie et al., 2022b; Chen et al., 2024; Arakawa et al., 2024; Yang et al., 2024b).
We implement SensorChat in two system variants to accommodate various user needs: SensorChatC (“C” stands for “Cloud”) for deployment on a cloud server and SensorChatE (“E” stands for “Edge”) for deployment on an edge device. SensorChatC leverages OpenAI’s GPT APIs (Ye et al., 2023; Achiam et al., 2023) for better interpretation of user queries, enabling real-time QA interactions on a server equipped with an NVIDIA A100 GPU (a10, 2025). On the edge side, SensorChatE operates entirely on an edge device, preserving user privacy albeit at the cost of increased latency. We test SensorChatE on an NVIDIA Jetson Orin NX module (jet, 2025).
We conduct comprehensive evaluations of SensorChat on a state-of-the-art QA dataset, mainly covering the quantitative queries. The dataset is carefully chosen to include long-term multimodal sensor data and practical QA samples aligned with users’ interests. Additionally, we recruited eight volunteers for a real-world user study, approved by the Institutional Review Board (IRB). Participants were instructed to carry a smartphone while following their daily routines for one to three days. At the end of the sensor data collection period, they were encouraged to interact with SensorChat, ask any questions they had, and provide their feedback.
@@
  • We propose SensorChat, the first end-to-end system for answering qualitative and quantitative questions in long-term monitoring using high-dimensional time series sensors. SensorChat is designed with a novel three stage pipeline, including question decomposition, sensor data query and answer assembly.
  • We introduce a novel contrastive sensor-text pretraining loss for training the sensor and text encoders offline, enabling accurate sensor embedding queries given arbitrary partial text during the online sensor data query.
  • We enhance the LLMs in SensorChat through in-context learning and fine-tuning, incorporating carefully designed prompts to improve the quality of question decomposition and answer assembly.
  • Comprehensive evaluations show that SensorChat improves answer accuracy by up to 26% on quantitative questions, compared to the best-performing state-of-the-art QA systems. A real-world user study confirms SensorChat ’s ability to answer qualitative and open-ended questions to users’ satisfaction. In terms of latency, SensorChatC enables real-time user interactions from a cloud server, while SensorChatE operates fully on local edge devices with higher latency.
@@ 2. Related Work
Question Answering using Sensor Data The QA problem has been extensively studied across various domains, including text (Rogers et al., 2023), visual (Schwenk et al., 2022), medical (Pal et al., 2022), and remote sensing (Hu et al., 2023). In the sensor domain, early works AI Therapist (Nie et al., 2022a) and its successor CaiTI (Nie et al., 2024) utilized smart home devices, such as Amazon Echo, to engage in conversations with users and assess mental well-being. DeepSQA (Xing et al., 2021) was the first to benchmark time-series sensor-based QA for human activity recognition. It introduced SQA-GEN, an automated QA generation tool that gathers 1-minute sensor readings and generates valid Q&A pairs by exhaustively searching six pre-defined question templates. They mainly focused on quantitative questions including time query, counting and action compare. The authors also evaluated traditional neural network models, including CNNs and LSTMs, finding that the ConvLSTM network with Compositional Attention achieved the highest QA accuracy.
Recent contributions have pioneered the integration of LLMs for generating more intuitive answers and better interpreting sensor data. Englhardt et al. (Englhardt et al., 2024), Health-LLM (Kim et al., 2024), and DrHouse (Yang et al., 2024b) converted physiological data from wearable devices, such as heart rates and daily step counts, into text prompts for LLMs, enabling more sophisticated medical and healthcare diagnoses. The latest Sensor2Text (Chen et al., 2024) and PrISM-Q&A (Arakawa et al., 2024) explored natural language interactions between users and wearable devices to understand and support daily activities using sensor data, such as advising on ”What should I do next with this?” Both approaches utilized LLMs as their backbone and fed sensor embeddings into the models.
While promising, existing systems exhibit significant limitations in handling long duration and complex sensor data required for accurate answers. Table 1 highlights the key differences between SensorChat and existing QA systems. In summary, SensorChat greatly enhances capabilities of existing systems by (1) answering questions based on long-duration sensor data spanning weeks or months, compared to the short windows of seconds or minutes in previous systems (Xing et al., 2021; Chen et al., 2024; Arakawa et al., 2024), and (2) encoding high-dimensional time-series data to extract fine-grained activity details for reasoning, unlike prior systems that are limited to low-dimensional sensor data (Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024b).
LLMs for Multimodal Reasoning Recent works investigated multimodal LLMs that transform other data modalities into a sequence of tokens for LLM integration (Zhang et al., 2023; Lin et al., 2024b). IMU2CLIP (Moon et al., 2023b) and TENT (Zhou et al., 2023) employed contrastive pretraining to align text with various timeseries sensor signals. Recent designs like AnyMAL (Moon et al., 2023a) and OneLLM (Han et al., 2024) proposed fine-tuning multimodal LLMs to process up to eight different modalities, including IMU time series, for reasoning tasks. However, direct integration of time series sensor data to LLMs is constrained by LLM’s inherent weakness in handling long-context inputs (Li et al., 2024; Gu and Dao, 2023), making them unsuitable for processing long-duration sensor signals. In contrast, SensorChat overcomes this limitation by incorporating a dedicated sensor data query stage, enabling it to handle long-term queries effectively.
Table 1. Comparing SensorChat and existing QA systems.
Existing QA systems for sensor data Long-duration High-dimensional
@@
DeepSQA (Xing et al., 2021), Sensor2Text (Chen et al., 2024), PrISM-Q&A (Arakawa et al., 2024) $\times$
Englhardt et al. (Englhardt et al., 2024), Health-LLM (Kim et al., 2024), DrHouse (Yang et al., 2024b) $\times$
@@
Mobile Systems for Daily Life Monitoring Researchers have explored a variety of sensing technologies for monitoring human lives, including built-in sensors on smart phones (Zhang et al., 2020), cameras (Radu and Henne, 2019), Wi-Fi signals (Yang et al., 2024a) and mmWave radar (Weng et al., 2024). Although these efforts have resulted in numerous open-source datasets (Reyes-Ortiz et al., 2012; Banos et al., 2014; Roggen et al., 2012; Vaizman et al., 2017) and powerful machine learning models (Ma et al., 2019; Xu et al., 2021; Deldari et al., 2022; Ouyang et al., 2022, 2023; Xu et al., 2023b), they fail to handle queries in natural language, which are more creative and open-ended. In this paper, we design SensorChat to facilitate long-term QA interactions based on time series sensor data. The methodology of SensorChat can be further integrated with other sensing modalities and applications.

### v2

@@ Abstract.
Natural language interaction with sensing systems is crucial for addressing users’ personal concerns and providing health-related insights into their daily lives. When a user asks a question, the system automatically analyzes the full history of sensor data, extracts relevant information, and generate an appropriate response. However, existing systems are limited to short-duration (e.g., one minute) or low-frequency (e.g., daily step count) sensor data. In addition, they struggle with quantitative questions that require precise numerical answers.
In this work, we introduce SensorChat, the first end-to-end QA system designed for daily life monitoring using long-duration, high-frequency time series data. Given raw sensor signals spanning multiple days and a user-defined natural language question, SensorChat generates semantically meaningful responses that directly address users’ concerns. SensorChat effectively handles both quantitative questions that require numerical precision and qualitative questions that require high-level reasoning to infer subjective insights. To achieve this, SensorChat uses an innovative three-stage pipeline including question decomposition, sensor data query, and answer assembly. The first and third stages leverage Large Language Models (LLMs) to interpret human queries and generate responses. The intermediate querying stage extracts relevant information from the complete sensor data history, which is then combined with the original query in the final stage to produce accurate and meaningful answers. Real-world implementation demonstrate SensorChat’s capability for real-time interactions on a cloud server while also being able to run entirely on edge platforms after quantization. Comprehensive QA evaluations show that SensorChat achieves up to 93% higher answer accuracy than state-of-the-art systems on quantitative questions. Additionally, a user study with eight volunteers highlights SensorChat’s effectiveness in answering qualitative and open-ended questions.
@@ 1. Introduction
In recent years, the rapid growth of Internet of Things (IoT) devices has made them a ubiquitous part of daily life, from smartphones and smart home accessories to wearable technologies (Choi et al., 2021). These devices, equipped with multimodal on-device sensors, enable continuous monitoring of human activities and environmental conditions. The large volumes of long-term sensor data collected by these devices can provide a comprehensive view of users’ daily routines and valuable insights into their health. While traditional research has focused on building classification models for tasks like human activity recognition (Ouyang et al., 2022, 2023; Xu et al., 2023b, a; Cao et al., 2024; Weng et al., 2024), recent work has highlighted the value of natural language interfaces for making sensor data more accessible and useful to humans (Xing et al., 2021; Nie et al., 2022; Yang et al., 2024; Arakawa et al., 2024; Ji et al., 2024b; Chen et al., 2024). Question Answering (QA) systems, where users ask questions and receive responses from the sensor systems, offer a flexible and intuitive way to engage with sensor data. Unlike passive lifelogs or visualizations (Aizawa et al., 2004; Takahashi et al., 2013; Hwang and Cho, 2014; Adiga et al., 2020; Xu et al., 2024b), QA enables users to actively express their information needs, making the system more responsive to their personalized concerns.
A useful QA system should be able to handle long-term sensor data and a wide range of practical queries. We envision users collecting sensor data through everyday mobile devices (like smartphones and smartwatches) equipped with multimodal sensors, such as IMUs, microphones, GPS, and compasses. These sensors operate continuously in the background, capturing rich streams of data that can be stored locally or sent to the cloud. When users have specific questions or want to seek insights, they interact with the QA system, which automatically understands their intent, analyzes historical raw sensor data to extract relevant information, and generates meaningful answers. The QA system can run either on mobile devices or in the cloud, depending on resource needs and privacy requirements. To build a truly practical QA system, we focus on long-duration sensor data that spans more than a day, enabling the analysis of extended or recurring behavioral patterns (e.g., daily routines or lifestyle trends) that are not observable in shorter time frames. In addition, the system should be capable of processing high-frequency raw time series data to capture fine-grained activities, like gym workouts, which cannot be detected through sparse sensor readings such as daily step counts. On the query side, existing research for daily life monitoring has primarily focused on two broad categories: quantitative and qualitative questions (Xing et al., 2021; Reichman et al., 2025; Henrich and Herschbach, 2000). Quantitative questions require precise, objective answers derived from sensor data (Xing et al., 2021; Reichman et al., 2025), often involving content such as true/false, location, activity, count, or time duration (e.g., “How long did I exercise yesterday?”). In contrast, qualitative questions involve interpreting sensor data to infer high-level, subjective insights (Henrich and Herschbach, 2000). Qualitative questions can be vague, open-ended and often require external knowledge. For example, answering “Did I exercise enough for a young adult?” requires understanding of the standard activity levels. We recognize that addressing both categories of questions is important in building a useful QA system. To clarify our problem scope, we summarize the key concepts in Table 1.
In this paper, we explore the core research question: How can we design a practical QA system that delivers accurate and meaningful answers to both quantitative and qualitative user queries, based on long-duration, high-frequency sensor data? We recognize three key challenges for this problem:
  1. (1)
    Processing long-duration and high-frequency sensor data is inherently challenging due to the massive volume of raw data generated. For instance, a continuously running IMU sensor with 9 axes at a 40Hz sampling rate can produce roughly 125MB of data per day. In practice, mobile devices are equipped with a wide array of sensors beyond IMUs, further increasing the complexity and size of the data. Extracting relevant information from such vast, multimodal time series data becomes a difficult task that involves pattern recognition and temporal query retrieval.
  2. (2)
    Effectively fusing variable duration of sensor data with textual questions poses a significant challenge to answer certain questions. For instance, a specific quantitative question like ”What was I doing at 10:00 AM yesterday?” requires the system to precisely locate and interpret sensor data from a specific timestamp. In contrast, a high-level qualitative question such as ”How would you rate my lifestyle?” demands a broader understanding of patterns across extended periods. Effectively addressing these diverse queries requires flexible reasoning across multiple time scales, along with the ability to align temporal sensor data with natural language queries.
  3. (3)
    Generating accurate and semantically meaningful answers makes QA fundamentally more challenging than traditional classification or journaling tasks. In contrast to conventional human activity recognition (Ouyang et al., 2022, 2023; Xu et al., 2023b, a; Cao et al., 2024; Weng et al., 2024), which typically involves selecting from 10–20 predefined activity classes, QA requires choosing the next appropriate token from a vocabulary of over 32,000 possibilities at each step (Touvron et al., 2023a). Unlike passive journaling systems (Aizawa et al., 2004; Takahashi et al., 2013; Hwang and Cho, 2014; Adiga et al., 2020; Xu et al., 2024b) that simply log events, QA systems need to understand the user’s intent and deliver relevant, satisfying responses. The response needs to be a numerically precise answer to a quantitative question or a thoughtful, semantically rich interpretation for a qualitative question. This high level of complexity demands deeper reasoning, language understanding, and tight alignment between sensor data and user queries.
Table 1. Key concepts that defines the problem scope of this paper. These concepts are motivated by practical user needs.
Concept Meaning
Long-duration data Sensor data spans more than a full day of user activity
High-frequency data Collected from high-frequency sensors (e.g., IMU raw signals), as opposed to sparse data like daily step counts or statistics (e.g., mean value)
Quantitative questions Require precise and objective answers based on sensor data, such as true/false, location, activity, count, or time duration
Qualitative questions Require high-level reasoning to infer subjective insights, such as work-life balance and social interactions, possibly involving external knowledge
Existing research falls short in addressing all above challenges. Current QA systems are limited to short sensor durations (Xing et al., 2021; Chen et al., 2024; Arakawa et al., 2024; Moon et al., 2023b; Han et al., 2024) (e.g., one minute, as shown in Fig. 1(a)) or low-frequency readings (Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024) (e.g., daily step counts and temperature, as shown in Fig. 1(b)). In Section 3.2, we demonstrate that directly applying these systems to long-duration, high-frequency sensor data results in poor answer quality due to the increased data volume and complexity. Although recent advances in Large Language Models (LLMs) and multimodal LLMs offer new ways to understand raw sensor data, current approaches are ineffective for handling long-duration, high-frequency sensor data in our setting. These approaches typically involve converting raw sensor signals into text (Ji et al., 2024a; Xu et al., 2024a; Hota et al., 2024; Liu et al., 2023), visualizing them as images (Yoon et al., 2024), or encoding them into embeddings before feeding them into an LLM (Moon et al., 2023a; Han et al., 2024; Mo et al., 2024; Chen et al., 2024). However, as detailed in Appendix A, uncompressed text or embeddings often exceed the token limit of LLMs (e.g., 8,092 tokens for GPT-4 (Achiam et al., 2023)), while compressing high-frequency sensor readings over a day into a single image results in loss of interpretability. To the best of our knowledge, no existing system effectively addresses all three challenges in our setting.
(a) Example QA in DeepSQA (Xing et al., 2021) for human activities. The sensor duration is limited to 1 minute per conversation.
(b) Example QA in DrHouse (Yang et al., 2024) for clinical diagnosis. The sensor data is limited to sparse sensor readings thus may overlook fine-grained activities.
(c) Example QA in SensorChat (our system). In contrast to DeepSQA (Xing et al., 2021) and DrHouse (Yang et al., 2024), SensorChat handles long-duration and high-frequency raw sensor signals to accurately answer quantitative questions (left) and offer insightful qualitative analyses of well-being (right). DeepSQA (Xing et al., 2021), when directly adapted to our sensor input, produces incorrect answers. The correct answer is in green while the incorrect one is in red.
Figure 1. A visual comparison existing QA systems for sensor data and SensorChat.
@@
In this paper, we introduce SensorChat, the first end-to-end system designed for answering both qualitative and quantitative queries in long-term sensor monitoring, using multimodal and high-frequency time series sensors. In contrast to prior QA systems (Xing et al., 2021; Chen et al., 2024; Arakawa et al., 2024; Moon et al., 2023b; Han et al., 2024; Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024), SensorChat is capable of processing long-duration and high-frequency sensor data spanning weeks or even months. SensorChat generates accurate answers to quantitative questions and semantically meaningful answers to qualitative questions, all based on the extracted sensor information. Figure 1(c) illustrates two QA examples using SensorChat. The input consists of raw multimodal sensor data collected over three days from a smartphone and a smartwatch. For clarity, we visualize the inferred user activity labels, with the x-axis showing time of day (00:00 AM to 11:59 PM) and the y-axis indicating separate days. Different activities are represented by color-coded bars. When asked a quantitative query like “What day did I spend the longest eating last week?”, SensorChat automatically understands the question and searches the full sensor history to return a precise answer. SensorChat can also handle open-ended qualitative questions, such as queries “How would you rate my lifestyle?” by reasoning on top of a detailed analysis of daily routines.
To achieve the goal, SensorChat introduces a novel three-stage design including question decomposition, sensor data query and answer assembly. The key design philosophy of SensorChat is to use LLMs in the first and third stages to correctly understand user questions and generate natural language answers, while inserting an explicit query stage in between for accurate sensor data processing and effective sensor-question fusion. SensorChat tackles the first challenge of processing long-duration, high-frequency sensor data by pretraining a sensor encoder and a text encoder using a novel contrastive loss. This pretraining enables the model to associate raw sensor signals with semantically meaningful text embeddings. Then, in the second stage, SensorChat uses a custom query procedure to identify relevant sensor segments from the full monitoring history based on the textual input, bridging sensor data and natural language. Lastly, to ensure high-quality responses, SensorChat uses in-context learning (Shao et al., 2023) and Chain-of-Thought (CoT) prompting (Chu et al., 2024) during question decomposition to correctly interpret user intent. In the answer assembly stage, SensorChat applies Low Rank Adaptation (LoRA) finetuning (Hu et al., 2021) to align answers with user preferences, based on a state-of-the-art dataset (Reichman et al., 2025). The three stages of SensorChat work collaboratively to deliver accurate and meaningful responses to both quantitative and qualitative questions over long-term, high-frequency multimodal sensor data.
We implement SensorChat in two system variants to accommodate various user needs: SensorChatC (for cloud deployment) and SensorChatE (for edge deployment). SensorChatC uses OpenAI’s GPT APIs (Ye et al., 2023; Achiam et al., 2023) to interpret user queries, offering real-time QA interactions on a server with an NVIDIA A100 GPU (a10, 2025). On the other hand, SensorChatE uses quantized LLMs and runs entirely on an edge device, preserving user privacy albeit at the cost of higher latency. We test SensorChatE on an NVIDIA Jetson Orin NX module (jet, 2025). We evaluate SensorChat through two sets of experiments to test different capabilities of the system. First, we test SensorChat on SensorQA (Reichman et al., 2025), a state-of-the-art QA dataset focused on quantitative questions. Next, we conduct a real-world user study with eight volunteers, approved by the Institutional Review Board (IRB), to assess SensorChat ’s performance on both qualitative and quantitative questions. This study also tests SensorChat’s generalizability to personalized mobile devices.
@@
  • We propose SensorChat, the first end-to-end system for answering qualitative and quantitative questions in long-term monitoring using high-frequency time series sensors. SensorChat is designed with a novel three stage pipeline, including question decomposition, sensor data query and answer assembly.
  • We introduce a novel contrastive sensor-text pretraining loss for training the sensor and text encoders offline. We also design a custom query procedure to ensure accurate sensor embedding queries given arbitrary text during the online question-answering.
  • We enhance the LLMs in SensorChat through in-context learning, CoT and fine-tuning, incorporating carefully designed prompts to improve the quality of question decomposition and answer assembly.
  • Evaluations on a quantitative QA dataset show that SensorChat improves answer accuracy by up to 93% compared to the best performing state-of-the-art QA systems. A real-world user study confirms that SensorChat provides meaningful and satisfying answers to users’ questions, including both quantitative and qualitative ones, while generalizing to personal mobile devices.
  • We implement SensorChat in two system variants, SensorChatC and SensorChatE. SensorChatC enables real-time user interactions from a cloud server, while SensorChatE operates fully on local edge devices with reasonable answer latency.
@@ 2. Related Work
Table 2. Comparison of SensorChat with existing sensor-based QA systems. SensorChat is the first QA system to provide accurate, meaningful answers based on long-duration, high-frequency time series sensor data.
Existing QA systems Long-duration High-frequency
@@
DeepSQA (Xing et al., 2021), Sensor2Text (Chen et al., 2024), PrISM-Q&A (Arakawa et al., 2024), OneLLM (Han et al., 2024) $\times$
Englhardt et al. (Englhardt et al., 2024), Health-LLM (Kim et al., 2024), DrHouse (Yang et al., 2024) $\times$
@@
Question Answering using Sensor Data In the sensing domain, prior QA systems can be broadly categorized based on the type and duration of sensor data used, as summarized in Table 2. One line of research focuses on understanding users’ current activities using real-time sensor readings, as shown in Fig. 1(a). For each QA instance, DeepSQA (Xing et al., 2021) processes one-minute IMU signals, Sensor2Text (Chen et al., 2024) uses 16-second body tracking data, PrISM-Q&A (Arakawa et al., 2024) analyzes short-segment smartwatch signals for procedural guidance via a voice assistant, OneLLM (Han et al., 2024) allows 10-second IMU signals as one of the eight modailities. All of these works are limited to short-duration sensor data, whereas directly applying these methods to long-duration sensor data in our setting leads to ineffective fusion between sensor and text, as we detail in Sec. 3.2. On the contrary, SensorChat supports understanding and reasoning over long-term data spanning more than a day. The second line of research uses long-term but sparse sensor data to support health-related queries and clinical assessments. Englhardt et al. (Englhardt et al., 2024), Health-LLM (Kim et al., 2024), and DrHouse (Yang et al., 2024) converted physiological signals (e.g., heart rate, step counts) into text prompts for LLMs, as exemplified in Fig. 1(b). However, these methods rely on low-frequency data, thus cannot capture fine-grained activities like gym sessions. In contrast, SensorChat processes raw, high-frequency sensor signals, capturing detailed user schedules and activity durations.
Multimodal LLMs using Sensor Data The rise of LLMs has sparked growing interest in integrating sensor data into LLMs for high-level reasoning. HARGPT (Ji et al., 2024a), Penetrative AI (Xu et al., 2024a), Hota et al. (Hota et al., 2024), and Liu et al. (Liu et al., 2023) are among the first works to convert raw sensor signals into text prompts for LLMs. By-My-Eyes (Yoon et al., 2024) selects visualization tools to present sensor data as visual prompts. Other efforts in multimodal LLMs (Moon et al., 2023a; Han et al., 2023) train adapters to convert raw sensor signals into embeddings, which are then used for finetuning LLMs. However, these methods struggle when tasked with interpreting the long-duration, high-frequency data in our scenario, as discussed in Appendix A. More recent approaches like IMU2CLIP (Moon et al., 2023b) and LLMSense (Ouyang and Srivastava, 2024) use classifiers to map raw signals into narrative logs for LLM input, but still suffer from LLMs’ limitations in handling long context (Li et al., 2024; Gu and Dao, 2023), i.e., input with more than 2K tokens. In contrast, SensorChat overcomes this limitation by incorporating a dedicated sensor data query stage, enabling it to handle long-term sensor queries effectively.